OpenStudy (anonymous):

Distributive Property! Simplify expression : -6(a+8)

OpenStudy (anonymous):

-6a-48

OpenStudy (lgbasallote):

let me show you an *example* of how to use distributive property \[\large 2(x + y + 1) \implies (2\times x) + (2\times y) + (2\times 1)\] \[\large \implies 2x + 2y + 2\]

OpenStudy (anonymous):

and if you're trying to solve for a then move -6a to the other side and then proceed to isolate a